# How Transliterations Improve Crosslingual Alignment

### v1

@@ 4.4 Models
Model-2. These models are trained on the concatenation of the original and transliterated data only with MLM. For example, we concatenate Data and Data and use the resulted data as the training data for pol-ukr pair.
Model-3. The training data is the same as the data used for Model-2. However, both MLM and TCM objectives are used in training. The final loss is the sum of MLM and TCM.
Model-4. The training data is the same as the data used for Model-2. However, both MLM and TLM objectives are used in training. The final loss is the sum of MLM and TLM.
Model-5. The training data is the same as the data used for Model-2. However, all objectives are used in training: MLM, TCM, and TLM. The final loss is the sum of MLM, TCM, and TLM.
@@ 4.5 Evaluation
Datasets and metric. Since sentence retrieval directly evaluates the quality of crosslingual alignment, we focus on the sentence retrieval task as our primary evaluation. We consider two datasets: SR-B and SR-F. SR-B contains 450 parallel sentences from the Bible in each language’s original script. SR-F contains 1,012 parallel sentences from Flores200 (Team, 2024), also in each language’s original script. We report top-1, top-5, and top-10 accuracy for each direction in each language pair.
Results and discussion. Results are reported in Table 2. We observe that Model-1 already achieves very good retrieval performance, suggesting that models can implicitly learn good crosslingual alignment even without any supervision signals, consistent with previous research findings (Pires et al., 2019; Dufter and Schütze, 2020). Surprisingly, Model-2 generally performs worse than Model-1, indicating that simply adding transliterations to the training data does not improve crosslingual alignment between the two languages in their original scripts. However, as long as any auxiliary learning objective is incorporated, retrieval performance increases. The TCM objective is particularly effective: Model-3 and Model-5 achieve the best overall retrieval performance across datasets for the pol-ukr and hin-urd pairs, respectively. Our findings can be summarized as follows: (1) vanilla MLM on related languages can already achieve good crosslingual alignment, (2) adding transliterations alone does not improve alignment, and (3) alignment is improved when any auxiliary objective is included, especially TCM, which directly operates on sentence-level representations.
@@ 5 Analysis
We interpret the results of §4.5 by establishing a connection between crosslingual alignment and different types of similarities (§5.2). We also analyze the dynamics of these similarities during the pretraining phase (§5.3). Finally, we provide insights on how crosslingual alignment influences downstream crosslingual transfer (§5.4). Our analysis primarily focuses on SR-B, as the impact of each component is more pronounced (Table 2). See Appendix §B for SR-F.
@@ 5.1 Defining Similarities
For a sentence written in its original script in language , we denote its transliteration in Latin script, its translation in language , and the transliterated translation. We then define the transliteration similarity as , the translation similarity as , the transliteration-translation similarity as , and the transliteration-transliteration similarity as , where takes a text as input and encodes it as a fixed-size representation. We mean-pool the output from the 8th layer to form such fixed-size representations. For simplicity, is always ukr (resp. urd) and is always pol (resp. hin) for the pol-ukr (resp. hin-urd) pair, as both and are the same when interchanging the languages. See Appendix §C for the other direction.
@@ 5.2 Similarities and Alignment
As discussed in §3, good alignment does not necessarily require a model to assign high similarity to matched pairs (translations). Instead, the model should be able to differentiate matched pairs from non-matched pairs. We display the translation similarity between matched sentence pairs and between random sentence pairs in Figure 1. We also compare the four types of similarities in each model in Figure 2.
@@
Augmentation alone improves similarity but not alignment. As shown in §4.5, simply adding transliterations to the training data does not improve crosslingual alignment. However, in Figure 2, we observe that translation similarity in Model-2 improves compared to Model-1 (from 77 to 80 in terms of the average similarity scores for the pol-ukr pair). This suggests that the increased lexical overlap in the added transliterated data (cf. Table 1) implicitly improves overall similarity. Unfortunately, for this model, the similarity between random sentence pairs is also increased, as shown in Figure 1, which is detrimental to alignment. This observation agrees with previous studies showing that encoder-only models can mistakenly assign high cosine similarity scores to both matched and random word pairs (Ethayarajh, 2019; Zhao et al., 2021).
@@
Auxiliary learning objectives improve alignments. Figure 1 shows that Model-4 improves the translation similarity compared to Model-2, thanks to the inclusion of TLM. Although the similarity between random pairs also increases, the similarity gap between matched pairs and random pairs is slightly enlarged, contributing to a modest improvement in retrieval performance and crosslingual alignment. TCM is even more effective than TLM at improving overall similarities, while simultaneously improving the discrimination between matched and random pairs (see Figure 2). This can be attributed to the contrastive objective, which not only encourages representations of paired sentences to be similar but also teaches the model to differentiate unpaired sentences. Consequently, we observe the highest alignment in Model-3 and Model-5 for both language pairs.
@@
Alignment can be improved even with a “bad" tokenizer. Unlike the pol-ukr pair, where the tokenizer already contains many Latin subwords due to Polish using the Latin script, the hin-urd pair does not use this script at all. Therefore, the tokenization results for the transliteration of Hindi or Urdu texts are “bad," often producing very long sequences composed of individual characters like “a" and “b". As a result, the overall transliteration-transliteration similarity is very high for all model variants, as shown in Figure 2, especially when no transliterated data is incorporated (Model-1). However, despite this “bad" tokenizer, the TCM approach significantly improves crosslingual alignment. This indicates that TCM does not necessarily rely on high-quality tokenizations of transliterated texts. In other words, its effectiveness is robust.
@@ 5.3 Similarity Dynamics during Pretraining
To analyze the “dynamics” of similarities, i.e., their variation during pretraining, we plot all four types of similarities for each model and each language pair at every 2K epochs in Figure 3.
@@
Increased lexical overlap improves the initial transliteration-transliteration similarity. Since transliterated sentences from two languages in a language pair share much more lexical overlap compared to sentences in their original scripts, we observe high transliteration-transliteration similarity and low translation similarity at the early stages of the training, as shown in Figure 3 (this effect is particularly clear in Model-1). This is even more so for the Hindi-Urdu pair, where the tokenizer often produces sequences of individual characters for any pair of transliterated sentences, resulting in extremely high transliteration-transliteration similarity. However, when transliterated data is not used in the training, this similarity gradually decreases in Model-1 for both language pairs.
Transliterations serve as an intermediary in improving translation similarity. When translated data is included, translation similarity increases more rapidly (as seen when comparing the similarity progression of other models with Model-1 in Figure 3). As all other types of similarities gradually decrease in Model-1, we can infer that the faster improvement in translation similarity is due to the inclusion of transliterated data, improving both transliteration similarity and transliteration-transliteration similarity. For example, the similarity between hin_Deva and hin_Latn is improved, so is the similarity between urd_Arab and urd_Latn. The improved lexical overlap also boosts the similarity between hin_Latn and urd_Latn. The combined effect ultimately leads to further improvement in the similarity between hin_Deva and urd_Arab. This effect is amplified when TCM is applied, as it directly optimizes the model for higher transliteration similarity.
5.4 Downstream Crosslingual Performance
Although better alignment is expected to help crosslingual transfer, Wu and Dredze (2020), Gaschi et al. (2023) show that better token-level alignment does not always improve performance. We aim to further explore the connection between the sentence-level alignment – our focus – and downstream crosslingual performance. We evaluate this connection using three datasets: SIB200 (Adelani et al., 2024) for text classification, WikiANN (Pan et al., 2017) for NER, and Universal Dependencies (de Marneffe et al., 2021) for Part-of-speech tagging. We report both in-language and crosslingual transfer performance for each language pair, with the results presented in Table 3.
@@ 5.4 Downstream Crosslingual Performance
Auxiliary objectives can be detrimental for in-language evaluation but beneficial for transfer. We observe a decrease in performance when training and evaluating on the same language. E.g., for SIB200, when training on Hindi and evaluating on Hindi, Model-3, 4 and 5 are worse than Model-1. Similar trends are observed for Polish and Urdu. Conversely, when training on one language and evaluating on another, there is often a performance improvement. This suggests that auxiliary objectives may negatively impact the quality of the representations within a specific language, resulting in worse in-language performance. However, better alignment enhances the similarity of representations for similar sentences across languages, which can be beneficial for zero-shot crosslingual transfer.
@@
Better crosslingual alignment does not always improve transfer. Although Model-3, 4, and 5 demonstrate better alignment compared to Model-1, the crosslingual transfer performance does not improve substantially—especially when considering the magnitude of alignment improvement seen in Table 2. This is particularly clear in sequential tasks like NER and POS, where all models achieve comparable performance, regardless of whether the evaluation is in-language or crosslingual. Even in SIB200, the improvement in Model-3, 4 and 5 is inconsistent: there is much better transfer performance for the directions and but slightly worse performance for the directions and . Therefore, our results suggest that better sentence-level crosslingual alignment does not consistently lead to improved crosslingual transfer. It should be noted that better word alignment (ground-truth word-level aligned data is required to evaluate this) may have a positive effect on the transfer of NER / POS.
@@ 6 Conclusion
We show that adding transliterated data can improve crosslingual alignment as transliterations act as an intermediary. This effect is particularly pronounced when auxiliary alignment objectives are applied, allowing models to better distinguish matched pairs from random pairs, thereby improving alignment. However, the improved alignment doesn’t consistently produce better downstream performance. Our theoretical and empirical analysis suggests that more research is needed to better understand the relationship between crosslingual alignment and crosslingual transfer.
@@
This work presents the first attempt to explain why the transliteration-augmented methods can improve crosslingual alignment, which usually requires parallel data in the training or fine-tuning. One possible limitation is the number of language pairs we consider: we use two language pairs where each pair contains two related languages that use different scripts. A possible future direction to expand this work is to explore more language pairs, or even involve more than two related languages in the training to investigate the effect of transliteration in a highly multilingual context.
Another possible limitation is that we focus on the sentence-level crosslingual alignment in this paper. For token-level crosslingual alignment, one can use (round-trip) word alignment can be used for evaluation, which has shown to be hard to find (ImaniGooghari et al., 2023).
@@ Appendix A Training Details
To pretrain different model variants for each language pair, we use the AdamW optimizer (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) with and . The initial learning rate is set to 5e-5. The effective batch size is 1,024 in each training step, with gradient accumulation set to 16 and 8 training instances (each instance actually contains a pair of sentences, see paragraph below for explanation) are used for each of the 8 Nvidia RTX 2080Ti GPUs (). We use FP16 training with mixed precision (Micikevicius et al., 2018). We store checkpoints every 2K steps and apply early stopping based on the best average performance on SR-B retrieval. The pretraining takes around 2 days for each model.
Except for Model-1, all other model variants double the training data because of including transliterated data. This can result in a different number of parameter update steps in an epoch between Model-1 and other models (the hyperparameters used by the AdamW optimizer will be different for each step), adding confounder variables to our analysis. To solve this problem, every instance in each batch is a pair of sentences in the pretraining. For Model-1, two identical sentences (in the original script) are used to form the pair, whereas the sentence (in the original script) and its transliteration are used to form the pair in other model variants. This setup ensures the total training steps in an epoch are the same for the pretraining of each model variant.
@@ Appendix B Additional Analysis on SR-F
We show the transliteration similarity between matched sentence pairs and between random sentence pairs from SR-F in Figure 4. We see a similar trend as we observe in SR-B. However, because each sentence in Flores Team (2024) is relatively simple and quite different from the other sentences in the dataset, the similarity gap between matched pairs and random pairs is already quite large. Therefore, the effect of including transliterations in the training data or additionally with transliteration-alignment objectives is marginal.
Similarly, we visualize the four types of similarities in each model for SR-F in Figure 5. The trend remains almost the same as for SR-B: including the transliterated data improves all similarities and the usage of additional alignment objectives can further improve different similarities.
We plot all four types of similarities measured using SR-F for each model and language pair for every 2K epochs in Figure 6. The trend is almost identical to what we observe when measuring the similarity using SR-B: including transliterations has a direct effect on transliteration-transliteration similarity and transliterations can implicitly improve the translation similarity since transliterations work as an intermediary.

### v2

@@ 4.3 Training Objectives
Transliteration Language Modeling (TLM).
This objective, proposed by Xhelili et al. (2024), is similar to the translation language modeling of Conneau and Lample (2019), where we use transliterations, instead of translations, to build sentence pairs in the objective. Following (Xhelili et al., 2024), we concatenate a sentence and its transliteration and perform MLM on the combined text. To predict a token masked in the original sentence, the model can either attend to tokens in the original script or their transliterations, and vice versa.
4.4 Models
We train a SentencePiece Unigram tokenizer (Kudo, 2018; Kudo and Richardson, 2018) on Data and Data for each language pair, respectively. We set the size of vocabularies to 30K for each pair. The tokenizers are not adapted to the transliterated data, i.e., Data and Data, in order to replicate the settings used by Liu et al. (2024b) and Xhelili et al. (2024), as they achieve surprisingly good performance without any tokenizer adaptation. As shown in Table 1, lexical overlap increases drastically for transliterated data even without learning subwords from it. We then train five model variants from scratch for each language pair to thoroughly explore the effect of each component of the transliteration-augmented pretraining. We introduce the 5 model variants as follows (training details are reported in §A).333MLM is used in each model variant because it is important for language modeling. Training the models from scratch only with TCM or TLM can result in bad language modeling ability and therefore such options are not considered in our study.
Model-1.
These models are trained on either Data or Data only with MLM.
Model-2.
These models are trained on the concatenation of the original and transliterated data only with MLM. For example, we concatenate Data and Data and use the resulted data as the training data for pol-ukr pair.
Model-3.
The training data is the same as the data used for Model-2. However, both MLM and TCM objectives are used in training. The final loss is the sum of MLM and TCM.
Model-4.
The training data is the same as the data used for Model-2. However, both MLM and TLM objectives are used in training. The final loss is the sum of MLM and TLM.
Model-5.
The training data is the same as the data used for Model-2. However, all objectives are used in training: MLM, TCM, and TLM. The final loss is the sum of MLM, TCM, and TLM.
@@ 4.5 Evaluation
Datasets and metric.
Since sentence retrieval directly evaluates the quality of crosslingual alignment, we focus on the sentence retrieval task as our primary evaluation. We consider two datasets: SR-B and SR-F. SR-B contains 450 parallel sentences from the Bible in each language’s original script. SR-F contains 1,012 parallel sentences from Flores200 (Team, 2024), also in each language’s original script. We report top-1, top-5, and top-10 accuracy for each direction in each language pair.
Results and discussion.
Results are reported in Table 2. We observe that Model-1 already achieves very good retrieval performance, suggesting that models can implicitly learn good crosslingual alignment even without any supervision signals, consistent with previous research findings (Pires et al., 2019; Dufter and Schütze, 2020). Surprisingly, Model-2 generally performs worse than Model-1, indicating that simply adding transliterations to the training data does not improve crosslingual alignment between the two languages in their original scripts. However, as long as any auxiliary learning objective is incorporated, retrieval performance increases. The TCM objective is particularly effective: Model-3 and Model-5 achieve the best overall retrieval performance across datasets for both pol-ukr and hin-urd pairs. The TLM objective is less effective compared with TCM but still helps to improve the alignment: Model-4 achieves worse performance than Model-3 but outperforms Model-1 and Model-2 in general. Our findings can be summarized as follows: (1) vanilla MLM on related languages with different scripts can already achieve good crosslingual alignment, (2) adding transliterated data in pretraining alone has a negative impact on crosslingual alignment, and (3) alignment is improved when any auxiliary objective is included, especially TCM, which directly operates on sentence-level representations.
@@ 5 Analysis
We interpret the results of §4.5 by establishing a connection between crosslingual alignment and four different types of similarities (§5.2). We also analyze the dynamics of these similarities during the pretraining phase (§5.3). Finally, we provide insights on how crosslingual alignment influences zero-shot crosslingual transfer performance in downstream tasks (§5.4). Our analysis in the following primarily focuses on SR-B, as the impact of each component is more pronounced (Table 2). See Appendix §B for additional analysis on SR-F.
@@ 5.1 Defining Similarities
For a sentence written in its original script in language , we denote as its transliteration in Latin script, as its translation in language , and as its transliterated translation. We then define the transliteration similarity as , the translation similarity as , the transliteration-translation similarity as , and the transliteration-transliteration similarity as , where takes a text as input and encodes it as a fixed-size representation. We mean-pool the output from the 8th layer to form such fixed-size representations. For simplicity, is always ukr (resp. urd) and is always pol (resp. hin) for the pol-ukr (resp. hin-urd) pair, as both and are the same when interchanging the languages. See Appendix §C for the other direction.
@@ 5.2 Similarities and Alignment
As discussed in §3, good crosslingual alignment does not necessarily require a model to assign high similarity to matched sentence pairs (translations). Instead, the model should be able to differentiate matched pairs from non-matched pairs to achieve better alignment. We display the similarity between matched sentence pairs and between random sentence pairs in Figure 1. We also compare the four types of similarities in each model in Figure 2.
Adding transliterated data alone improves similarity but not alignment.
@@
As shown in §4.5, simply adding transliterated data to the training data does not improve crosslingual alignment. However, in Figure 2, we observe that translation similarity in Model-2 improves compared to Model-1 (from 77 to 80 in terms of the average similarity scores for the pol-ukr pair). This suggests that the increased lexical overlap in the added transliterated data (cf. Table 1) implicitly improves overall similarities. Unfortunately, for this model, the similarity between random sentence pairs is also increased, as shown in Figure 1, which is detrimental to crosslingual alignments. This observation agrees with some previous studies showing that encoder-only models can mistakenly assign high cosine similarity scores to both matched and random word pairs (Ethayarajh, 2019; Zhao et al., 2021).
Auxiliary learning objectives improve alignments.
@@
Figures 1 and 2 show that Model-4 improves the translation similarity – similarity between the matched pairs, compared to Model-2, thanks to the inclusion of TLM. Although the similarity between random pairs also increases, the similarity gap between matched pairs and random pairs is slightly enlarged, contributing to a modest improvement in retrieval performance and crosslingual alignment. TCM is even more effective than TLM at improving overall similarities (see Figure 2), while simultaneously improving the gap between matched and random pairs (see Figure 1). This can be attributed to the contrastive objective, which not only encourages representations of paired sentences to be similar but also teaches the model to differentiate unpaired sentences. Consequently, we observe the best alignments in Model-3 and Model-5 for both language pairs.
Alignment can be improved even with a “bad” tokenizer.
@@
Unlike the pol-ukr pair, where the tokenizer already contains many Latin-script subwords due to Polish using the Latin script, the hin-urd pair does not use this script at all. Therefore, the tokenization results for the transliteration of Hindi or Urdu texts are “bad”: the tokenizer often produces very long sequences composed of individual characters like “a” and “b”. As a result, the overall transliteration-transliteration similarity is very high for all model variants, as shown in Figure 2, especially when no transliterated data is incorporated in the pretraining data (Model-1). However, despite such a “bad” tokenizer, the TCM objective significantly improves crosslingual alignment. This indicates that TCM does not necessarily rely on high-quality tokenizations of transliterated texts. In other words, its effectiveness is robust.
@@ 5.3 Similarity Dynamics During Pretraining
To analyze the dynamics of similarities, i.e., their variation during pretraining progression, we plot all four types of similarities for each model and each language pair at every 2K steps in Figure 3.444Each step corresponds to a single update of the parameters during pretraining.
Similarities involving transliteration decrease if no transliterated data is added.
We observe high transliteration-transliteration similarity at the early stages of the pretraining in Model-1. However, because no transliterated data is added, this similarity, along with other similarities involving transliteration, gradually drops, as shown in Figure 3. When transliterated data is included, all transliteration-related similarities increase throughout pretraining (this effect is particularly clear when comparing Model-2 with Model-1). This trend can be explained by the fact that language in its original script and in the Latin script are intrinsically the same language. The model can quickly learn alignment between them as long as the transliterated data is included, even if explicit alignment objectives, e.g., TCM, are not used.
Transliterations serve as an intermediary in improving translation similarity.
When transliterated data is included, translation similarity increases more rapidly (as seen when comparing the similarity progression of other models with Model-1 in Figure 3). As all other similarities gradually decrease in Model-1, we can infer that the faster improvement in translation similarity shown in other models is due to the improved transliteration similarity and transliteration-transliteration similarity, by including transliterated data. For example, the similarity between hin_Deva and hin_Latn (transliteration similarity) is improved, so is the similarity between urd_Arab and urd_Latn (transliteration similarity if we refer to urd as the source language and the trend can be seen in Figure 9). The improved lexical overlap in transliterated data boosts the similarity between hin_Latn and urd_Latn. The combined effect ultimately leads to further improvement in the similarity between hin_Deva and urd_Arab (translation-translation similarity). We can observe this intermediary effect is amplified when TCM is applied, as it directly optimizes the model for higher transliteration similarity (this similarity is much higher in Model-3 and Model-5 compared to other models).
5.4 Downstream Crosslingual Performance
Although better alignment is expected to help crosslingual transfer, Wu and Dredze (2020), Gaschi et al. (2023) show that better token-level alignment does not always improve performance. We aim to further explore the connection between the sentence-level alignment – our focus – and downstream crosslingual performance. We evaluate this connection using three datasets: SIB200 (Adelani et al., 2024) for text classification, WikiANN (Pan et al., 2017) for named entity recognition (NER) NER, and Universal Dependencies (de Marneffe et al., 2021) for Part-of-speech tagging (POS). We report both in-language and crosslingual transfer performance for each language pair, with the results presented in Table 3.
@@ 5.4 Downstream Crosslingual Performance
Auxiliary objectives can be detrimental for in-language evaluation but beneficial for transfer.
@@
We observe a decrease in performance when training and evaluating on the same language. For instance, for SIB200, when training on Hindi and evaluating on Hindi, Model-3, 4 and 5 are worse than Model-1. Similar trends are observed for Polish and Urdu. Conversely, when training on one language and evaluating on another, there is often a performance improvement. This suggests that auxiliary objectives may negatively impact the quality of the representations within a specific language, resulting in worse in-language performance. However, better alignment enhances the similarity of representations for similar sentences across languages, which can be beneficial for zero-shot crosslingual transfer.
Better crosslingual alignment does not always improve transfer.
@@
Although Model-3, 4, and 5 demonstrate better alignment compared to Model-1 (cf. Table 2), the crosslingual transfer performance does not substantially improve – especially when considering the magnitude of alignment improvement seen in Table 2. This is particularly clear in sequential tasks like NER and POS, where all models achieve comparable performance, regardless of whether the evaluation is in-language or crosslingual. Even for SIB200, the improvement in Model-3, 4 and 5 is inconsistent: there is much better transfer performance for the directions and but slightly worse performance for the directions and . Therefore, our results suggest that better sentence-level crosslingual alignments do not consistently lead to improved crosslingual transfer, especially for sequential tasks such as NER and POS. We conjecture that the lack of explicit token-level alignment objectives with word-level aligned data in our models might explain why we do not see improvements in these tasks, similar to the findings from Chaudhary et al. (2020) and Xhelili et al. (2024).
@@ 6 Conclusion
Our work presents the first in-depth study exploring why and how transliterations contribute to better crosslingual alignment. We show that adding transliterated data can improve crosslingual alignment as transliteration acts as an intermediary between pairs of mutual translations. This effect is particularly pronounced when auxiliary alignment objectives are applied, allowing models to better distinguish matched pairs from random pairs, thereby improving the overall alignment. However, our empirical results also show that improved alignment does not consistently produce better downstream performance, suggesting more research is needed to better understand the relationship between crosslingual alignment and crosslingual transfer.
7 Future Work
We see possible future work to overcome the limitations mentioned in the Limitations Section. A possible direction to expand this work is to explore more language pairs, or even involve more than two related languages in the training to investigate the effect of transliteration in a highly multilingual context. For further assessing token-level alignment, one possible way is to use word-level aligned data, which is unfortunately not much in the community. As an alternative, one can use (round-trip) word alignments, which have been shown to be very hard to find (ImaniGooghari et al., 2023).
@@ 7 Future Work
This work presents the first attempt to explain why the transliteration-augmented methods can improve crosslingual alignment, which usually requires parallel data in the training or fine-tuning. One possible limitation is the number of language pairs we consider: we only use two language pairs, each of which contains two related languages that use different scripts. Another possible limitation is that we only focus on the sentence-level crosslingual alignment in this paper and do not discuss the token-level crosslingual alignment.
@@ Appendix A Training Details
To pretrain different model variants for each language pair, we use the AdamW optimizer (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) with and . The initial learning rate is set to 5e-5. The effective batch size is 1,024 in each training step, with gradient accumulation set to 16 and 8 training instances (each instance contains a pair of sentences, see paragraph below for explanation) are used for each of the 8 NVIDIA RTX 2080Ti GPUs (). We use FP16 training with mixed precision (Micikevicius et al., 2018). We store checkpoints every 2K steps and apply early stopping based on the best average performance on SR-B retrieval task. The pretraining takes around 2 days for each model.
Except for Model-1, all other models double the training data due to the inclusion of transliterated data. This can result in a different number of parameter updates in an epoch between Model-1 and other models (the hyperparameters used by the AdamW optimizer will be different for each step), adding confounding variables to our analysis. To solve this problem, every instance in each batch is a pair of sentences in the pretraining. For Model-1, two identical sentences (in the original script) are used to form a pair, whereas a sentence and its transliteration are used to form a pair in other models. This setup ensures the total training steps in an epoch are the same for all models.
@@ Appendix B Additional Analysis on SR-F
We show the similarity between matched sentence pairs and between random sentence pairs from SR-F in Figure 4. We see a similar trend as for SR-B. However, because each sentence in Flores Team (2024) is relatively simple and quite different from the other sentences in the dataset, the similarity gap between matched and random pairs is already quite large. Therefore, the effect of including transliterations or auxiliary objectives is marginal. Similarly, we visualize the four types of similarities in each model for SR-F in Figure 5. The trend remains almost the same as for SR-B: including the transliterated data improves all similarities and the usage of transliteration-based alignment objectives can further improve overall similarities.
We plot all four types of similarities measured using SR-F for each model and language pair throughtout pretraining in Figure 6. The trend is almost identical to what we observe when measuring the similarity using SR-B: including transliterations has a direct effect on transliteration-transliteration similarity and transliterations can implicitly improve the translation similarity since transliterations work as an intermediary.